OpenStudy (anonymous):

If TRS=126 and the area of the smaller sector is 35pi , calculate the area of R.

OpenStudy (anonymous):

Someone please help...i really need this

OpenStudy (lalaly):

area of the sector = pi r^2 · θ/360 35π = pi r^2 · 126/360 35 = 0.35r^2 r = 10 area of circle = pi r^2= 100pi

OpenStudy (anonymous):

haha yup...right again...more questions coming right up

OpenStudy (anonymous):

thanks!!

OpenStudy (lalaly):

;)np